A delivery truck travels 21 blocks north, 16 blocks east, and 26 blocks south. What is its final displacement from the origin? Assume the blocks are equal length.
1
Expert's answer
2018-02-06T08:13:06-0500
X-coordinate is
16.
Y-coordinate is
21-26=-5.
The magnitude of displacement is
d=√(〖16〗^2+(-5)^2 )=17 blocks.
The direction is
θ=tan^(-1)⁡〖5/16〗=17° east of south.